LBTO Technical/D/Service Night

OSA: D.G.Huerta

ISp: J.Morris

Manager: K.Newton

ISA: B.Rothberg

======================================

CLOSED FOR SNOW

UT 01:30 - Testing LUCI-2 calibrations for spectroscopy

UT 03:10 - Running 0.5" G210 K@2.2 cals for UGC6 data 0050-0069 flats

UT 03:24 - Cal unit now stuck trying to stabilize voltage

UT 03:28 - Everytime a lamp change is made (on/off or different type of lamp) the voltage is reset for the grating.

UT 03:57 - Talking to Alexander, trying to reinitizlize the insturment to keep it from being stuck in loop of stable/unstable

UT 04:03 - Alexander is restarting LUCI2

UT 04:04 - not sure why it keeps showing 1.00micron G210 with clear and K filters, when script gave it 2.2micron CWL

UT 04:08 - Alexander restarted LUCI2 and restarted 0.5" G210@2.2 cal script, 0070 first exposure

UT 04:14 - script caused error. all that was suppoesd to happen was a lamp turns off, yet LUCI-2 resets the voltage setting and the whole instrument manager shows an error

UT 04:19 - Running H-band script, 210, 0.5" 1.65CWL. Now it shows unstable 0080 should be first exposure of set

UT 04:24 - Failed on H-band - flipped to 1.0 micron in display

UT 04:30 - Shut everything down. LUCI-2 in safe position, sent an email describing issue to luci email list.

Done for the night!

-- %USERSIG{BarryRothberg - 2015-10-30}%

Comments


 
Topic revision: r2 - 30 Oct 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback